OpenStudy (anonymous):

Find f(x)·g(x) if f(x) = 3x and g(x) = x + 1. A. 3x2 + 3x B. 3x + 1 C. 4x + 1 D. 3x2 + 1

OpenStudy (anonymous):

is it B ?

mathslover (mathslover):

No. sorry @Savannah2D0pe you have done somewhere wrong.

mathslover (mathslover):

See f(x) . g(x) = 3x(x+1) right?

OpenStudy (anonymous):

yeah .

OpenStudy (anonymous):

so 4x+1 ?

mathslover (mathslover):

No : a(b+c ) = \(ab + ac\) so \(3x ( x+1 ) = 3x (x ) + 3x(1) \) ^ That is called distributive property

OpenStudy (anonymous):

i just dont get it ...

mathslover (mathslover):

See : \(a ( b + c ) = ab + ac\) that is 3x(4x+1) = 3x*4x + 3x*!

mathslover (mathslover):

@Savannah2D0pe http://teachers.henrico.k12.va.us/math/ms/c20708/01NumberSense/1-5DistributiveProp.html ^ That would surely help you understand

mathslover (mathslover):